OpenStudy (mertsj):

@wasiqss How is it 50 pi radians per second? If you contradict an answer you should provide a reason.

OpenStudy (anonymous):

original question?

OpenStudy (wasiqss):

the speed was 25 m/s,so it means 25 rev in one second, hence 1 rev = 2pie so multiply 2 pie with 25

OpenStudy (mertsj):

25 m per second does not mean 25 rev in 1 sec since 1 rev is pi m not 1 meter

OpenStudy (turingtest):

I wanna know the question as well now

OpenStudy (anonymous):

it would help if we know size of wheel

OpenStudy (mertsj):

a truck is moving at a rate of 90km/hr, and the diameter of its wheel is 1 meter. find the angular speed of the wheels in radians per second.

OpenStudy (turingtest):

it is a circumference of pi

OpenStudy (mertsj):

Of course. That is what I said.

OpenStudy (turingtest):

I know, just backing you up

OpenStudy (anonymous):

It is 50 rad/sec

OpenStudy (wasiqss):

diameter was one meter

OpenStudy (mertsj):

Yep so 1 revolution is C = Pi times diameter or pi meters

OpenStudy (anonymous):

90000 ------- = cycle per hours 2 pi .5 90000 ------- 1/3600 = cycle per second 2 pi .5 90000 ------- *2pi*1/3600 =(90000/.5)*1/3600= radian per second

OpenStudy (wasiqss):

i rev= 2pie, hence 25 rev means 50 pie

OpenStudy (anonymous):

who said 25 rev?